The average expenditure on Valentine's Day was expected to be $100.89 (USA Today, February 13, 2006). Do male and female consumers differ in the amounts they spend? The average expenditure in a sample survey of 50 male consumers was $132.48, and the average expenditure in a sample survey of 34 female consumers was $65.27. Based on past surveys, the standard deviation for male consumers is assumed to be $32, and the standard deviation for female consumers is assumed to be $16. The z value is 2.576. a. What is the point estimate of the difference between the population mean expenditure for males and the population mean expenditure for females? b. At 99% confidence, what is the margin of error? c. Develop a 99% confidence interval for the difference between the two population means. .... to ....
Answer #1

a)

point estimate of the difference between the population mean expenditure for males and the population mean expenditure for females = difference in sample means = x̅1 - x̅2 =    132.48   -   65.27   = 67.21

b)

sample #1   ------->   male
mean of sample 1,    x̅1=   132.48
population std dev of sample 1,   σ1 =    32
size of sample 1,    n1=   50
      
sample #2   --------->   female
mean of sample 2,    x̅2=   65.27
population std dev of sample 2,   σ2 =    16
size of sample 2,    n2=   34

std error , SE =    √(σ1²/n1+σ2²/n2) =    5.2924          
margin of error, E = Z*SE =    2.576   *   5.292   =   13.6323

c)
                  
difference of means = x̅1 - x̅2 =    132.48   -   65.27   =   67.210
confidence interval is                   
Interval Lower Limit= (x̅1 - x̅2) - E =    67.210   -   13.632   =   53.5777
Interval Upper Limit= (x̅1 - x̅2) + E =    67.210   +   13.632   =   80.8423
a 99% confidence interval for the difference between the two population means ( 53.5777 , 80.8423)